A capacitor charged from a 50V DC supply is discharged across a charge-measuring instrument and found to have carried a charge of 10uC. What is the capacitance of the capacitor and how much energy was stored in it?
1
Expert's answer
2021-02-11T17:11:22-0500

The capacitance of the capacitor is:


"C=\frac{Q} {U}=\frac{10\cdot10^{-6}} {50} =0.2\space uF"


The stored energy is:


"W=\frac{QU} {2} =\frac{10\cdot 10^{-6}\cdot 50} {2} =0.25 \space mJ."